# AgentTTS: Large Language Model Agent for Test-time Compute-optimal Scaling Strategy in Complex Tasks

### v1

@@ Abstract
Test-time scaling (TTS) enhances the performance of large language models (LLMs) by allocating additional compute resources during inference. However, existing research primarily investigates TTS in single-stage tasks; while many real-world problems are multi-stage complex tasks, composed of a sequence of heterogeneous subtasks with each subtask requires LLM of specific capability. Therefore, we study a novel problem: the test-time compute-optimal scaling in multi-stage complex tasks, aiming to select suitable models and allocate budgets per subtask to maximize overall performance. TTS in multi-stage tasks introduces two fundamental challenges: (i) The combinatorial search space of model and budget allocations, combined with the high cost of inference, makes brute-force search impractical. (ii) The optimal model and budget allocations across subtasks are interdependent, increasing the complexity of the compute-optimal search. To address this gap, we conduct extensive pilot experiments on four tasks across six datasets, deriving three empirical insights characterizing the behavior of LLMs in multi-stage complex tasks. Informed by these insights, we propose AgentTTS, an LLM-agent-based framework that autonomously searches for compute-optimal allocations through iterative feedback-driven interactions with the execution environment. Experimental results demonstrate that AgentTTS significantly outperforms traditional and other LLM-based baselines in search efficiency, and shows improved robustness to varying training set sizes and enhanced interpretability.
@@ 1 Introduction
Test-time scaling (TTS), which refers to allocating additional computational resources during inference, has shown promising results in improving the performance of large language models (LLMs) (Brown et al., 2024; Wu et al., 2025; Zeng et al., 2025; Snell et al., 2025; Liu et al., 2025a; Yue et al., 2025; Muennighoff et al., 2025). For example, Brown et al. (2024) scaled inference compute via repeated sampling of candidate solutions, using a verifier to select the best prediction, enabling a weak small model to outperform single-sample state-of-the-art models. Despite its effectiveness, existing methods are primarily designed for single-stage tasks, where the model performs a single function, such as mathematical problem solving (Wu et al., 2025; Brown et al., 2024; Zeng et al., 2025; Snell et al., 2025; Liu et al., 2025a; Muennighoff et al., 2025; Jin et al., 2025) or code generation (Brown et al., 2024; Stroebl et al., 2024). However, many real-world applications involve multi-stage complex tasks, for which compute allocation remains underexplored. These tasks involve a sequential execution of heterogeneous subtasks to accomplish complex objectives. Representative examples include retrieval-then-generation QA systems (Lewis et al., 2020), waterfall-style software development (comprising requirement analysis, system design, coding, and testing) (Qian et al., 2024), and multi-agent task automation (such as task decomposition, tool selection, and parameter prediction) Shen et al. (2024). Each subtask within a multi-stage workflow often requires a model with specific capabilities (Jiang et al., 2023). For example, in a retrieval-then-generation QA task, retrieval benefits from large models with strong long-context understanding, while generation can achieve competitive performance using smaller models with repeated sampling (see Fig. 1(a)(b)). Therefore, multi-stage tasks demand models with diverse types and levels of capabilities, challenging existing TTS methods.
@@
To address these challenges, we first conduct preliminary experiments to characterize the behavior of LLMs under test-time scaling in multi-stage tasks. From experiments, we derive three generalizable insights across four task types: (1) Different subtasks exhibit distinct preferences between large and small models; (2) Increasing test-time compute initially improves performance, but beyond a certain point, additional compute yields diminishing or no gains; (3) The compute allocated to earlier subtasks influences the scaling dynamics and compute needs of downstream subtasks. Guided by these insights, we propose AgentTTS, a novel LLM-agent-based framework designed to efficiently search for compute-optimal budget allocations in multi-stage tasks. Recent works (Liu et al., 2024a; Zhang et al., 2024a; Yang et al., 2025b; Morris et al., 2024; Luo et al., 2024) have shown that LLM-based agents are effective in planning and searching for hyperparameter optimization. Given the capability of LLMs in understanding and following instructions, AgentTTS integrates our observed test-time scaling insights into the LLM-agent search process. AgentTTS consists of three key components: the Agent, Archive, and Environment. The Agent begins by generating an initial set of trials based on Insight 1, which guides model preferences for each subtask. These trials are executed by the Environment, which evaluates them on the actual task platform and returns performance feedback. The Archive stores a history of generated trials, guidelines, and feedback. In subsequent stages, the Agent produces new trials and guidelines informed by Insights 2 and 3. This iterative process continues until a predefined stopping criterion is met. By leveraging LLM-based search, AgentTTS offers two advantages: (i) interpretability, through explicit guideline generation that explains decision rationales; and (ii) robustness, in navigating non-smooth search spaces commonly found in test-time scaling.
@@ 2 Related Work
Test-time Scaling and Compute-optimal Strategy. Test-time scaling (TTS) enhances LLM performance by allocating additional compute during inference (Brown et al., 2024; Wu et al., 2025). Existing methods fall into two categories: sequential scaling(Madaan et al., 2023; Gou et al., 2024; Snell et al., 2025; Muennighoff et al., 2025; Yang et al., 2025a), which iteratively refines outputs but depends on good initial responses, and parallel scaling(Brown et al., 2024; Zeng et al., 2025; Stroebl et al., 2024; Sun et al., 2024; Gui et al., 2024; Snell et al., 2025; Liu et al., 2025a; Wu et al., 2025), which generates multiple outputs and aggregates them using reward-based selection (e.g., repeated sampling (Brown et al., 2024; Zeng et al., 2025), Best-of-N (Sun et al., 2024), or tree search (Wu et al., 2025)). Recent work reduces reliance on reward models by using LLMs as fusers (Jiang et al., 2023; Li et al., 2025; Saad-Falcon et al., 2024; Cao et al., 2025). Parallel scaling is preferred for complex tasks due to better scalability and broader solution coverage (Snell et al., 2025). Hence, we adopt repeated sampling with fusion. Research on test-time compute-optimal scaling shows that small models with optimal strategies can outperform larger ones (Brown et al., 2024; Wu et al., 2025; Liu et al., 2025a; Yue et al., 2025; Snell et al., 2025; Wang et al., 2024b). Approaches include difficulty-aware model selection (Snell et al., 2025), reward-guided voting (Wu et al., 2025), and budget-aware prompting (Yue et al., 2025). However, they focus on single-stage tasks, while we extend this to multi-stage tasks, where budgets must be adaptively allocated across interdependent subtasks to maximize overall performance. A more detailed discussion is given in Appendix A.11.1.
LLMs for Hyperparameter Optimization. LLMs have become powerful tools for hyperparameter optimization (HPO), surpassing traditional AutoML techniques such as Bayesian Optimization (BO) (Jones et al., 1998; Shahriari et al., 2015) by leveraging contextual reasoning and prior knowledge (Gu et al., 2024). LLM-based HPO research generally follows two directions: (1) reducing the search space and (2) directly generating hyperparameter configurations. For the former, LLMs have been used to prune large search spaces in NAS and HPO (Yu et al., 2023; Morris et al., 2024; Luo et al., 2024; Liu et al., 2024b), as seen in GPT-NAS (Yu et al., 2023), AutoM3L (Luo et al., 2024), and Llambo (Liu et al., 2024b). For the latter, recent works (Chen et al., 2023; Zheng et al., 2023; Zhang et al., 2023a; Achiam et al., 2023; Hollmann et al., 2023; Zhang et al., 2024a; Lin et al., 2025; Zhang et al., 2024b) treat LLMs as autonomous optimizers. Systems like AutoMMLab (Yang et al., 2025b), GENIUS (Zheng et al., 2023), MLCopilot (Zhang et al., 2024a), and AgentHPO (Liu et al., 2024a) refine trials via feedback and experience. We extend this line of work to compute-optimal test-time scaling in multi-stage tasks. A more detailed introduction of related work is given in Appendix A.11.2.
@@ 3 Preliminary Knowledge and Problem Definition
Test-time Scaling Mode: Repeated Sampling with Fusion. We adopt the Repeated Sampling with Fusion strategy for test-time scaling, as it does not rely on additional reward models or verifiers compared to Best-of-N (Sun et al., 2024; Gui et al., 2024) or tree search algorithms (Wu et al., 2025; Yu et al., 2024), and it offers greater scalability compared to sequential scaling (Zeng et al., 2025; Brown et al., 2024). Given a problem $p$, a language model $M$ with parameters $\theta$, test-time scaling is performed by increasing the number of repeated samples $k$. To stimulate diverse generations, we set the temperature hyperparameter to 0.9 throughout. A fusion prompt is then used to aggregate the multiple candidate solutions generated through repeated sampling:
@@ 4.1 Experimental Setting
Tasks, Datasets, and Models We conduct preliminary experiments on four multi-stage tasks: (i) Retrieval-based Question Answering using 2WikiMultiHopQA (Ho et al., 2020) and HotpotQA (Yang et al., 2018) datasets, (ii) Knowledge Graph Question Answering using CWQ (Talmor and Berant, 2018) and WebQSP (Yih et al., 2015), (iii) Task Automation, using TaskBench (Shen et al., 2024), and (iv) Automated Software Development, using ChatDev (Qian et al., 2024). Subtasks differ in prompt and generation lengths; retrieval uses longer prompts, while QA needs longer generations. We consider models ranging from 3B to 72B. Details of task specifications, datasets, models, and evaluation metrics are provided in Appendix A.4 and A.5.
@@
Compute Cost Metric and Budget Definition. Different stakeholders prioritize different cost metrics. Consumers are typically concerned with monetary expenses, such as the price per million input and output tokens. LLM researchers focus on computational cost, such as inference FLOPs, due to constraints of limited GPU memory. Discussion of using API price as a cost metric is in Appendix A.3. In this work, we adopt inference FLOPs as the primary cost metric throughout. The corresponding budget conversion is stated in the theory below. The proof is in Appendix A.2.
@@ 4.2 Preliminary Experimental Results and Insights
We use the FLOPs-based budget conversion in Eq. 4 to analyze performance variance with increasing sample count and compute budget, and to guide budget allocation in our method in the next section. Fig. 1 shows how subtask performance in retrieval-based question answering varies with increasing test-time sampling and computational budget (FLOPs). In Fig. 1 (a), performance on the retrieval subtask exhibits a strong positive correlation with model size, while smaller models provide only marginal improvements even with increased budget. This suggests that retrieval benefits from larger models, where fewer high-capacity samples yield better results. In contrast, Fig. 1 (b) shows that for the question answering subtask, under a limited budget (e.g., $<10^{14}$ FLOPs), LLaMA-3 3B and LLaMA-3 8B outperform LLaMA-3 70B, indicating a preference for smaller models. Similar trends are observed in Fig. 10 (a)(b) and across other benchmarks (see Fig. 11-14 in Appendix A.6). These discrepancies arise from subtask-specific demands: retrieval emphasizes long-context understanding, favoring large models; while question answering primarily involves extracting information from retrieved content, where smaller models excel and can better exploit test-time scaling via repeated sampling. Thus, different subtasks exhibit distinct preferences between language models and sampling frequency, depending on the specific capabilities required by each subtask (Insight 1).
@@
[…]
@@ 5 AgentTTS: Agent for Test-time Scaling Budget Allocation
where each trial is represented as $c=(M_{1},B_{1},\dots,M_{n},B_{n})$, $M_{i}\in\mathcal{M}_{i}$ is the selected model, and $B_{i}$ is its assigned budget for subtask $i$. $B$ denotes the total compute budget. $M_{i,\max}$ is the largest available model for subtask $i$, and the configuration $(...,M_{i},B_{i},...)$ indicates that other subtasks are fixed to their respective available largest models with one-pass inference.
@@ 6 Experiments
Experimental Setup. We conduct experiments across six datasets covering four distinct task categories, as detailed in Appendix A.4. Baselines include two groups of hyperparameter optimization: traditional machine learning methods and recent LLM-based methods. The traditional baselines include Bayesian optimization (BO) (Jones et al., 1998; Shahriari et al., 2015) and random search. For LLM-based approaches, we consider AgentHPO (Liu et al., 2024a), MLCopilot (Zhang et al., 2024a), and LLM_ZS. Given that these LLM-based methods were initially proposed for hyperparameter tuning, we adapt them to the test-time budget allocation. Both AgentHPO and MLCopilot employ a two-phase pipeline: (1) feedback-driven guideline generation and (2) guideline-informed candidate trial generation. They differ in the initialization of guidelines: AgentHPO generates initial guidelines autonomously via the LLM, while MLCopilot initializes guidelines based on performance trends observed in similar tasks. In contrast, LLM_ZS directly prompts the LLM to generate candidate trials without relying on guidelines or external references. Further details of each baseline are in Appendix A.10. All methods execute 50 iterations of search on a training set comprising 50 samples and are subsequently evaluated on a test set containing 500 samples. We use GPT-o3-mini (OpenAI, 2025) as the LLM search agent. The default budget is set as the sum of the minimum budget required for one-pass inference on each subtask using the largest model.
Main Results and Analysis. Fig. 3 shows the search trajectories of our method and the baselines for test-time compute-optimal budget allocation on the test set. The search time and test-set performance of the best trials are reported in Tab. 1. We observe: (i) Our method outperforms baselines in both search efficiency and test-set performance in Fig. 3, 15, and Tab. 1. First, ours achieves top results on most tasks and requires fewer trials, even when baselines eventually reach optimality. Second, our method reduces search time, confirming higher computational efficiency. These gains come from leveraging empirical insights to guide the search toward promising trials. Despite similar performance on the training set, our method often generalizes better, as shown in Tab. 1. For example, on 2WikiMultiHopQA, it exceeds the next best methods by 2% on the test set. This suggests that Insight 2 helps identify minimal budgets and avoid redundant sampling, improving generalization. (ii) Compared to LLM-agent-based AgentHPO and MLCopilot, our method converges faster, showing integrated insights improve search efficiency on compute-optimal allocation. (iii) Traditional methods such as BO often get stuck in local optima due to the non-smooth landscape, while random search is noise-tolerant but inefficient. In contrast, LLM-based methods use prior hyperparameter tuning knowledge to bypass suboptimal regions, achieving better search performance.
Figure 3: Performance trajectories across various search methods over 50 trials. X-axis: trial count; Y-axis: best performance up to each trial. The best score is obtained from the optimal trial in a prior grid search and serves as the benchmark for all methods. More results are in Appendix A.7.
@@
Figure 5: Interpreting AgentTTS Through Integrated Empirical Insights. 2Wiki means 2WikiMultiHopQA. A complete case is in Appendix A.8.
Interpretability of Budget Allocation To illustrate the interpretability benefit from integrating prior insights into the LLM agent for test-time compute allocation, we present three case studies on 2WikiMultiHopQA in Fig. 5, showing Insight 1 enables subtask-specific model use, favoring large models for retrieval and small ones for QA, to guide efficient trials.
@@
Performance under varying budget settings To assess the search efficiency under different budgets, we evaluate AgentTTS under two budget settings on 2WikiMultiHopQA: 500500500 (low budget: only one subtask can reach its optimum) and 200020002000 (high, full optima reachable but with larger search spaces). As shown in Fig. 6, we can conclude: (1) Our method finds optimal configurations under low budgets, confirming the benefit of integrated insights on trade-off subtask interdependence; (2) Across low to high budget settings, our method consistently outperforms baselines in search efficiency, indicating that the insight-informed LLM agent remains robust to the expanded search space.

### v2

@@ Abstract
Test-time scaling (TTS) enhances the performance of large language models (LLMs) by allocating additional compute resources during inference. However, existing research primarily investigates TTS in single-stage tasks; while many real-world problems are multi-stage complex tasks, composed of a sequence of heterogeneous subtasks with each subtask requires LLM of specific capability. Therefore, we study a novel problem: the test-time compute-optimal scaling in multi-stage complex tasks, aiming to select suitable models and allocate budgets per subtask to maximize overall performance. TTS in multi-stage tasks introduces two fundamental challenges: (i) The combinatorial search space of model and budget allocations, combined with the high cost of inference, makes brute-force search impractical. (ii) The optimal model and budget allocations across subtasks are interdependent, increasing the complexity of the compute-optimal search. To address this gap, we conduct extensive pilot experiments on four tasks across six datasets, deriving three empirical insights characterizing the behavior of LLMs in multi-stage complex tasks. Informed by these insights, we propose AgentTTS, an LLM-agent-based framework that autonomously searches for compute-optimal allocations through iterative feedback-driven interactions with the execution environment. Experimental results demonstrate that AgentTTS significantly outperforms traditional and other LLM-based baselines in search efficiency, and shows improved robustness to varying training set sizes and enhanced interpretability.  [^1] Code link: https://github.com/FairyFali/AgentTTS/
@@ 1 Introduction
Test-time scaling (TTS), which refers to allocating additional computational resources during inference, has shown promising results in improving the performance of large language models (LLMs) [2, 45, 55, 33, 19, 54, 26]. For example, Brown et al. [2] scaled inference compute via repeated sampling of candidate solutions, using a verifier to select the best prediction, enabling a weak small model to outperform single-sample state-of-the-art models. Despite its effectiveness, existing methods are primarily designed for single-stage tasks, where the model performs a single function, such as mathematical problem solving [45, 2, 55, 33, 19, 26, 14] or code generation [2, 34]. However, many real-world applications involve multi-stage complex tasks, for which compute allocation remains underexplored. These tasks involve a sequential execution of heterogeneous subtasks to accomplish complex objectives. Representative examples include retrieval-then-generation QA systems [16], waterfall-style software development (comprising requirement analysis, system design, coding, and testing) [29], and multi-agent task automation (such as task decomposition, tool selection, and parameter prediction) [32]. Each subtask within a multi-stage workflow often requires a model with specific capabilities [13]. For example, in a retrieval-then-generation QA task, retrieval benefits from large models with strong long-context understanding, while generation can achieve competitive performance using smaller models with repeated sampling (see Fig. 1(a)(b)). Therefore, multi-stage tasks demand models with diverse types and levels of capabilities, challenging existing TTS methods.
@@
To address these challenges, we first conduct preliminary experiments to characterize the behavior of LLMs under test-time scaling in multi-stage tasks. From experiments, we derive three generalizable insights across four task types: (1) Different subtasks exhibit distinct preferences between large and small models; (2) Increasing test-time compute initially improves performance, but beyond a certain point, additional compute yields diminishing or no gains; (3) The compute allocated to earlier subtasks influences the scaling dynamics and compute needs of downstream subtasks. Guided by these insights, we propose AgentTTS, a novel LLM-agent-based framework designed to efficiently search for compute-optimal budget allocations in multi-stage tasks. Recent works [20, 56, 49, 25, 23] have shown that LLM-based agents are effective in planning and searching for hyperparameter optimization. Given the capability of LLMs in understanding and following instructions, AgentTTS integrates our observed test-time scaling insights into the LLM-agent search process. AgentTTS consists of three key components: the Agent, Archive, and Environment. The Agent begins by generating an initial set of trials based on Insight 1, which guides model preferences for each subtask. These trials are executed by the Environment, which evaluates them on the actual task platform and returns performance feedback. The Archive stores a history of generated trials, guidelines, and feedback. In subsequent stages, the Agent produces new trials and guidelines informed by Insights 2 and 3. This iterative process continues until a predefined stopping criterion is met. By leveraging LLM-based search, AgentTTS offers two advantages: (i) interpretability, through explicit guideline generation that explains decision rationales; and (ii) robustness, in navigating non-smooth search spaces commonly found in test-time scaling.
@@ 2 Related Work
Test-time Scaling and Compute-optimal Strategy. Test-time scaling (TTS) enhances LLM performance by allocating additional compute during inference [2, 45, 39]. Existing methods fall into two categories: sequential scaling [24, 6, 33, 26, 48], which iteratively refines outputs but depends on good initial responses, and parallel scaling[2, 55, 34, 35, 9, 33, 19, 45], which generates multiple outputs and aggregates them using reward-based selection (e.g., repeated sampling [2, 55], Best-of-N [35], or tree search [45]). Recent work reduces reliance on reward models by using LLMs as fusers [13, 17, 30, 3]. Parallel scaling is preferred for complex tasks due to better scalability and broader solution coverage [33]. Hence, we adopt repeated sampling with fusion. Research on test-time compute-optimal scaling shows that small models with optimal strategies can outperform larger ones [2, 45, 19, 54, 33, 41, 40]. Approaches include difficulty-aware model selection [33], reward-guided voting [45], and budget-aware prompting [54]. However, they focus on single-stage tasks, while we extend this to multi-stage tasks, where budgets must be adaptively allocated across interdependent subtasks to maximize overall performance. A more detailed discussion is given in Appendix A.10.1.
LLMs for Hyperparameter Optimization. LLMs have become powerful tools for hyperparameter optimization (HPO), surpassing traditional AutoML techniques such as Bayesian Optimization (BO) [15, 31] by leveraging contextual reasoning and prior knowledge [8]. LLM-based HPO research generally follows two directions: (1) reducing the search space and (2) directly generating hyperparameter configurations. For the former, LLMs have been used to prune large search spaces in NAS and HPO [52, 25, 23, 21], as seen in GPT-NAS [52], AutoM3L [23], and Llambo [21]. For the latter, recent works [4, 62, 58, 1, 12, 56, 18, 57] treat LLMs as autonomous optimizers. Systems like AutoMMLab [49], GENIUS [62], MLCopilot [56], and AgentHPO [20] refine trials via feedback and experience. We extend this line of work to compute-optimal test-time scaling in multi-stage tasks. A more detailed introduction of related work is given in Appendix A.10.2.
@@ 3 Preliminary Knowledge and Problem Definition
Test-time Scaling Mode: Repeated Sampling with Fusion. We adopt the Repeated Sampling with Fusion strategy for test-time scaling, as it does not rely on additional reward models or verifiers compared to Best-of-N [35, 9] or tree search algorithms [45, 53], and it offers greater scalability compared to sequential scaling [55, 2]. Given a problem $p$, a language model $M$ with parameters $\theta$, test-time scaling is performed by increasing the number of repeated samples $k$. To stimulate diverse generations, we set the temperature hyperparameter to 0.9 throughout. A fusion prompt is then used to aggregate the multiple candidate solutions generated through repeated sampling:
@@ 4.1 Experimental Setting
Tasks, Datasets, and Models We conduct preliminary experiments on four multi-stage tasks: (i) Retrieval-based Question Answering using 2WikiMultiHopQA [11] and HotpotQA [50] datasets, (ii) Knowledge Graph Question Answering using CWQ [36] and WebQSP [51], (iii) Task Automation, using TaskBench [32], and (iv) Automated Software Development, using ChatDev [29]. Subtasks differ in prompt and generation lengths; retrieval uses longer prompts, while QA needs longer generations. We consider models ranging from 3B to 72B. Details of task specifications, datasets, models, and evaluation metrics are provided in Appendix A.4 and A.5.
@@
Compute Cost Metric and Budget Definition. Different stakeholders prioritize different cost metrics. Consumers are typically concerned with monetary expenses, such as the price per million input and output tokens. LLM researchers focus on computational cost, such as inference FLOPs, due to constraints of limited GPU memory. Discussion of using API price as a cost metric is in Appendix 6. In this work, we adopt inference FLOPs as the primary cost metric throughout. The corresponding budget conversion is stated in the theory below. The proof is in Appendix A.2.
@@ 4.2 Preliminary Experimental Results and Insights
We use the FLOPs-based budget conversion in Eq. 4 to analyze performance variance with increasing sample count and compute budget, and to guide budget allocation in our method in the next section. Fig. 1 shows how subtask performance in retrieval-based question answering varies with increasing test-time sampling and computational budget (FLOPs). In Fig. 1 (a), performance on the retrieval subtask exhibits a strong positive correlation with model size, while smaller models provide only marginal improvements even with increased budget. This suggests that retrieval benefits from larger models, where fewer high-capacity samples yield better results. In contrast, Fig. 1 (b) shows that for the question answering subtask, under a limited budget (e.g., $10^{14}$ FLOPs), LLaMA-3 3B and LLaMA-3 8B outperform LLaMA-3 70B, indicating a preference for smaller models. Similar trends are observed in Fig. 10 (a)(b) and across other benchmarks (see Fig. 11-14 in Appendix A.6). These discrepancies arise from subtask-specific demands: retrieval emphasizes long-context understanding, favoring large models; while question answering primarily involves extracting information from retrieved content, where smaller models excel and can better exploit test-time scaling via repeated sampling. Thus, different subtasks exhibit distinct preferences between language models and sampling frequency, depending on the specific capabilities required by each subtask (Insight 1).
@@
Fig. 1 (b-d) show how the performance of the question answering subtask varies under different levels of retrieval quality, with F1 scores approximately 0.80, 0.60, and 0.35, respectively. When high-quality retrieval is provided by a larger model (Fig. 1 (b)), we observe: (1) the optimal test-time budget for question answering is reached earlier. For instance, LLaMA-3 8B peaks at 10 samples (budget around $2\times 10^{13}$ FLOPs); and (2) smaller models (3B and 8B) may outperform the larger model (70B) under the same compute budget. In contrast, when retrieval quality is low (Fig. 1 (c)(d)), the peak performance is delayed. For example, with $\text{Ret\_F1}=0.6$, LLaMA-3 8B peaks at 20 samples (budget about $3\times 10^{13}$ FLOPs), while with $\text{Ret\_F1}=0.35$, even 90 samples (budget near $10^{14}$ FLOPs) still do not reach peak performance. Moreover, under low retrieval quality, smaller models are less likely to outperform larger ones. For instance, in Fig. 1 (d), the peak performance of the 3B and 8B models does not surpass even the lowest performance of the 70B model. These findings indicate that the scaling behavior of a subtask is affected by the performance of its preceding subtasks. Poor retrieval increases downstream task difficulty, requiring the model to compensate for missing information, where larger models are more beneficial. Similar trends are observed in Fig. 10 (b-d) and other benchmarks in Appendix A.6. We conclude that budget allocation for a preceding subtask impacts the model selection and optimal budget of subsequent subtasks (Insight 3).
@@ 5 AgentTTS: Agent for Test-time Scaling Budget Allocation
[…]
@@ 6 Experiments
Experimental Setup. We conduct experiments across six datasets covering four distinct task categories, as detailed in Appendix A.4. Baselines include two groups of hyperparameter optimization: traditional machine learning methods and recent LLM-based methods. The traditional baselines include Bayesian optimization (BO) [15, 31] and random search. For LLM-based approaches, we consider AgentHPO [20], MLCopilot [56], and LLM_ZS. Given that these LLM-based methods were initially proposed for hyperparameter tuning, we adapt them to the test-time budget allocation. Both AgentHPO and MLCopilot employ a two-phase pipeline: (1) feedback-driven guideline generation and (2) guideline-informed candidate trial generation. They differ in the initialization of guidelines: AgentHPO generates initial guidelines autonomously via the LLM, while MLCopilot initializes guidelines based on performance trends observed in similar tasks. In contrast, LLM_ZS directly prompts the LLM to generate candidate trials without relying on guidelines or external references. Further details of each baseline are in Appendix A.9. All methods execute 50 iterations of search on a training set comprising 50 samples and are subsequently evaluated on a test set containing 500 samples. We use GPT-o3-mini [27] as the LLM search agent. The default budget is set as the sum of the minimum budget required for one-pass inference on each subtask using the largest model.
@@
Main Results and Analysis. Fig. 3 shows the search trajectories of our method and the baselines for test-time compute-optimal budget allocation on the training set. The search time and test-set performance of the best trials are reported in Tab. 1. We observe: (i) Our method outperforms baselines in both search efficiency and test-set performance in Fig. 3 and Tab. 1. First, ours achieves top results on most tasks and requires fewer trials, even when baselines eventually reach optimality. Second, our method reduces search time, confirming higher computational efficiency. These gains come from leveraging empirical insights to guide the search toward promising trials. Despite similar performance on the training set, our method often generalizes better, as shown in Tab. 1. For example, on 2WikiMultiHopQA, it exceeds the next best methods by 2% on the test set. This suggests that Insight 2 helps identify minimal budgets and avoid redundant sampling, improving generalization. (ii) Compared to LLM-agent-based AgentHPO and MLCopilot, our method converges faster, showing integrated insights improve search efficiency on compute-optimal allocation. (iii) Traditional methods such as BO often get stuck in local optima due to the non-smooth landscape, while random search is noise-tolerant but inefficient. In contrast, LLM-based methods use prior hyperparameter tuning knowledge to bypass suboptimal regions, achieving better search performance.
@@
Interpretability of Budget Allocation To illustrate the interpretability benefit from integrating prior insights into the LLM agent for test-time compute allocation, we present three case studies on 2WikiMultiHopQA in Fig. 5. The first row shows Insight 1 enables subtask-specific model use, favoring large models for retrieval and small ones for QA, to guide efficient trials. The second row shows Insight 2 pinpoints the optimal sampling range (5–50), narrowing the search space. The third row shows Insight 3 promotes balanced allocation, favoring one-sample “high-capacity” retrieval and “lower-cost” QA configuration, enabling effective trade-offs for better overall performance.
@@
Performance under varying budget settings To assess the search efficiency under different budgets, we evaluate AgentTTS under two budget settings on 2WikiMultiHopQA: $500$ (low budget: only one subtask can reach its optimum) and $2000$ (high, full optima reachable but with larger search spaces). As shown in Fig. 6, we can conclude: (1) Our method finds optimal configurations under low budgets, confirming the benefit of integrated insights on trade-off subtask interdependence; (2) Across low to high budget settings, our method consistently outperforms baselines in search efficiency, indicating that the insight-informed LLM agent remains robust to the expanded search space.
Performance under varying temperatures To assess the impact of temperature on test-time scaling, we conduct additional experiments on the 2Wiki dataset. The retriever is fixed to Qwen2.5-72B (producing a single retrieval result), and we vary the number of samples and temperature values for the QA model LLaMA 3 3B. As shown in Tab. 2, higher temperatures (e.g., 0.9) improve performance with multiple samples, while lower temperatures (e.g., 0.1) are more effective for single-sample cases. This suggests that higher temperatures enhance output diversity and fusion in sampling-based reasoning, whereas lower temperatures promote more stable, deterministic outputs. These results align with prior findings [28], which show that temperature-induced diversity improves sample efficiency in reasoning tasks.
API price as the cost metric
Table 2: EM scores of LLaMA 3 3B at different temperature values on the 2Wiki dataset. The retriever is fixed to Qwen2.5-72B (single retrieval result).
Samples Temp = 0.1 Temp = 0.5 Temp = 0.9
1 0.70 0.64 0.66
10 0.70 0.78 0.78
20 0.72 0.76 0.78
30 0.74 0.78 0.72
40 0.74 0.76 0.78
50 0.74 0.72 0.78
60 0.74 0.78 0.80
70 0.76 0.72 0.74
80 0.76 0.76 0.78
90 0.74 0.72 0.76